OpenStudy (ashy98):

Can someone help me?(:

OpenStudy (ashy98):

Solve the following for y: 2x + 2y = 8 A.) y = −x + 4 <---- My answer B.) y = x + 4 C.) y = 3x D.) y = −3x

OpenStudy (anonymous):

looks good 2x+2y=8 2y=-2x+8 y=(-2x+8)/2 y=-x+4

OpenStudy (ashy98):

thank you & can you help me with some more I think i have done?(:

OpenStudy (anonymous):

yeah, sure

OpenStudy (ashy98):

The function p(x) = 2x +1 determines how many pizzas need to be purchased for an after school meeting, where x is the number of students at the meeting. The club leader uses r(p(x)) to find the amount of money to bring for the pizza purchase. The function r(x) = 4x − 3. Solve for how much money to bring when there are 9 students in the meeting. A.) 19 B.) 33 C.) 73 <----- D.) 84

OpenStudy (anonymous):

to find the amount of pizzas, we plug in p(9) we get p(9)=19 r(p(9))=r(19)=(4*19)-3=76-3=73 sorry that took a while, distracted rn